# Interaction induced reentrance of Bose glass and quench dynamics of Bose gases in twisted bilayer and quasicrystal optical lattices

### v1

@@ III.2 Reentrance between superfluid and Bose glass phases
In the previous section, we discussed the system at relatively low filling factors. From the particle density distribution in Fig. 1(b), we observe that the average number of particles per lattice site is less than 1, indicating the influence of the on-site interaction is very weak. This thus raises the question of whether different physical phenomena might arise when the system’s filling factor is higher. To explore this, we present the results for a system with a filling factor $\rho=0.08$, and the corresponding results are summarized in Fig. 2, which includes the phase diagram and typical real-space distributions. Compared to the phase diagram at low filling factors, we observe a reentrant transition at higher filling factors, as shown in Fig. 2(a). Focusing on the dotted line in this phase diagram, we see that as the on-site interaction $U$ increases, the system first transitions from the BG phase to the SF phase. However, as $U$ continues to increase, the system transitions back from the SF phase to the BG phase. This indicates that the on-site interaction can drive a reentrant transition to the BG phase.
@@
To understand this reentrant behavior, let us first check the typical real-space distributions of the system as $U$ changes are shown in Fig. 2(b). When the on-site interaction $U/J=1.25$ is relatively small compared to the external potential field strength $M_{r}/J=8$, the system consists of isolated SF clusters. As we gradually increase the on-site interaction to $U/J=1.75$, a percolated SF region (marked in blue) appears in the $S_{i}$ distribution shown in Fig. 2(b), indicating that the system has transitioned from the BG phase to the SF phase.
@@
The mechanism of this stage of transition is actually similar to the one of the lower filling case presented in Sec.III.1. Specifically, when the on-site interaction $U$ increases, the most significant interaction effect occurs at lattice sites where the average number of particles is greater than 1. At these sites, the particles repel each other, leading to a redistribution. As a result, SF paths emerge during this repulsion process, and the system transitions from the BG phase to the SF phase.
@@
However, as we further increase $U/J=3$, the system exhibits behavior different from that at low filling factors. As shown in the lower right plot of Fig. 2(b), the percolated SF region disappears from the distribution of $S_{i}$ at $U/J=3$, indicating the system transitions from the SF phase back to the BG phase again. To understand this stage of transition, we notice that although percolated SF regions can form in the SF phase, the increasing repulsive interaction strength can still cause the particle density distribution to undergo further rearrangement, causing percolated SF regions breaking up into isolated SF islands, thus driving the system entering the BG phase again. However, if particle density of each site is sufficiently low, this would indicate the local state of each site only consists of zero or single occupation state. As a consequence, further increasing on-site interaction strength can hardly impose any physical influence on the system, which is consistent with what we can observe in the last three columns of plot shown in Fig. 1(b) for the low filling case. In fact, at even high filling factors, one can observe even more times of the reentrance (see Appendix B for more numerical results at even higher filling factors).
@@
Figure 3: (a) Phase diagram of the system at $N=100$ with quench directions marked with arrows. (b) Time dependence of IPR of the corresponding quench dynamics. (c) Time dependence of percolation probability $\mathcal{P}$ of the corresponding quench dynamics. (d) Typical SF order parameter and discrete field distributions during different quench processes marked in (a). See text for more details.
@@ III.3 Quench dynamics
Thus far, we have focused on the static properties of the system. Now let us turn to the dynamical properties, in particular, here we shall focus on the quench dynamics of the system. Specifically, via time-dependent Gutzwiller approach (Schiró and Fabrizio, 2010; Bloch et al., 2008; Zakrzewski, 2005) (see Appendix C for technique details of the approach), we investigate the quench dynamics of two different types, namely, intra-phase quench and inter-phase quench.
For intra-phase quenches, where the system remains within the same phase (e.g., from the BG/SF phase to the BG/SF phase), we observe no qualitative changes in the system’s dynamics. As shown in Fig. 3, neither the time evolution of IPR nor the percolation probability $\mathcal{P}$ exhibit significant changes [see Figs. 3(b, c)]. Similarly, the typical SF order parameter and the discrete field distributions at different time instants remain largely unchanged [see Fig. 3(d)].
@@
For inter-phase quenches, where the system undergoes a transition between different phases (e.g., from SF to BG, or vice versa), we observe more significant qualitative changes. Specifically, for these quenches, the initial state corresponds to the ground state configuration at the two red dots in Fig. 3(a) with $(U/J=2,M_{r}/J=5)$ in the SF phase and $(U/J=2,M_{r}/J=10)$ in the BG phase. In the quench from the SF phase to the BG phase, the percolation probability $\mathcal{P}$ [illustrated by the green line in Fig. 3(c)] rapidly decreases from nearly 1 to 0, indicating the swift transition of the system from SF to BG. Once the system reaches the BG phase, $\mathcal{P}$ remains at 00, reflecting that the system remains in the BG phase. This behavior arises because the BG phase consists of isolated SF islands, which restrict particle movement to these disconnected regions. As a result, the IPR remains relatively constant with only minor fluctuations, corresponding to slight variations in the spatial distribution of particles over time [see the green line in Fig. 3(b)].
@@
For the quench from the BG phase to the SF phase, the percolation probability $\mathcal{P}$ fluctuates between 0 and non-zero values for times $tJ>5$, signaling the system’s transition between the BG and SF phases [see the blue line in Fig. 3(c)]. Unlike the quench from SF to BG, where particles are confined to isolated SF islands, in the BG-to-SF quench, particles can move freely between SF clusters, allowing the formation of SF paths. However, at certain moments, the system may lose its SF path and transition back to the BG phase. This phenomenon gradually fades over time, and as shown in Fig. 3, the periods during which $\mathcal{P}=0$ become shorter, eventually leading the system to settle into a steady SF phase. This trend is also reflected in the time evolution of the IPR in Fig. 3(b). Initially, in the BG phase, the particles are localized at specific lattice sites, resulting in a relatively large IPR. As the system evolves, particles diffuse through space, causing a gradual decrease in the IPR. Although the IPR may occasionally increase, it generally decreases over time, indicating that the system is eventually transitioning to the SF phase.
@@ IV CONCLUSIONS
Our work demonstrates that ultracold bosonic systems in optical lattices with an aperiodic external potential exhibit a rich interplay between disorder and interaction effects. We have shown that at low filling factors, an increase in the on-site repulsive interaction triggers a delocalization process that transforms a BG phase, characterized by isolated and disconnected SF regions, into a robust SF phase with a percolated network. This transition is clearly evidenced by the evolution of the percolation probability and IPR. At higher filling factors, we uncover a striking reentrant behavior wherein the system initially transitions from a BG to a SF phase as the interaction strength increases, but further enhancement of the interaction causes the SF network to fragment, reverting the system back to a BG phase. Moreover, our analysis of quench dynamics reveals markedly different transient behaviors for intra-phase and inter-phase quenches, highlighting the sensitive dependence of the system’s dynamical evolution on the initial state and interaction parameters. These findings not only deepen our understanding of localization phenomena in interacting Bose gases but also offer valuable insights for ongoing experiments utilizing twisted bilayer and quasicrystalline optical lattice platforms, where tuning disorder and interaction effects is pivotal for exploring novel many-body phases.
@@ Acknowledgements.
This work is supported by the NKRDPC (Grant No. 2022YFA1405304), NSFC (Grant Nos. 12275089 and 12004118), and the Guangdong Basic and Applied Basic Research Foundation (Grants Nos. 2024A1515010188, 2023A1515012800, and 2021A1515010212), and Guangdong Provincial Key Laboratory (Grant No. 2020B1212060066).
Appendix A TECHNICAL DETAILS ON CALCULATING PERCOLATION PROBABILITY FROM DISCRETE FILED
In this appendix, we provide technique details on calculating the percolation probability $\mathcal{P}$ from the discrete field $S_{i}$. The percolation probability $\mathcal{P}$ is defined in Reference (Johnstone et al., 2021) as
@@ Appendix A TECHNICAL DETAILS ON CALCULATING PERCOLATION PROBABILITY FROM DISCRETE FILED
where $N_{\textrm{span}}$ is the number of SF sites in a percolating cluster, and $N_{\varphi}$ is the total number of sites with finite $\varphi$. In our numerical calculations, $N_{\textrm{span}}$ and $N_{\varphi}$ are derived from the discrete field $S_{i}$ (Niederle and Rieger, 2013), which can be constructed by the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$ and SF order parameter distribution $\langle\hat{b}_{i}\rangle$ described in the main text. Specifically, we label sites with integer occupancy numbers as MI sites $\left(S_{i}=0\right)$ and sites with non-integer occupancy numbers as SF sites $\left(S_{i}=1\right)$. According to the definition of $S_{i}$, $N_{\varphi}$ corresponds to the total number of sites where $S_{i}=1$, i.e., $N_{\varphi}=\sum_{i}^{N_{\textrm{lat}}}S_{i}$. To calculate $N_{\textrm{span}}$, it is necessary to identify the percolating cluster. The criterion for percolation, as outlined in (Barman et al., 2013), is the existence of at least one SF path that spans the entire system, either from left to right or from top to bottom. Once such an SF path is identified, the SF cluster that contains this SF path is considered the percolating cluster. The total number of lattice sites in this percolating cluster is then defined as $N_{\textrm{span}}$. Clearly, as shown by the calculation of the percolation probability $\mathcal{P}$ described above, this quantity can be used to distinguish the SF phase from the BG phase. In the BG phase, there are no percolating SF clusters, which results in a percolation probability of 0. In contrast, in the SF phase, the presence of percolating SF clusters leads to a non-zero percolation probability.
@@ Appendix B NUMERICAL RESULTS AT EVEN HIGHER FILLING FACTORS
As discussed in the main text, we hypothesized that multiple reentrance phenomena could be observed in the system with an even higher particle filling. Indeed, as anticipated, this behavior is demonstrated in Fig. 4 below.
Figure 4: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $1024$ particles on a $51\times 51$ square lattice. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dotted line indicated in the phase diagram. (b) Configurations of the particle density distribution $\left\langle n_{i}\right\rangle$, SF order parameter distribution $\left\langle b_{i}\right\rangle$ and the discrete field $S_{i}$ of four points on the dotted line ($M_{r}/J=25$). See text for more details.
We plotted the phase diagram for a particle filling of 1024, which reveals two distinct peaks. Additionally, we included a series of real-space distribution diagrams showing the effect of varying the interaction strength $U$ on the lattice, with the changes in $U$ indicated by the horizontal lines in the phase diagram. As shown in the inset of the phase diagram, the percolation probability of the ground state exhibits multiple transitions between 0 and non-zero values, corresponding to the system’s behavior as it transitions from the BG phase to the SF phase, then back to the BG phase, and finally returning to the SF phase. The particle density distribution diagram Fig. 4(b) further illustrates that, as the particle filling increases, the system undergoes more significant changes due to the enhanced on-site interaction strength. Moreover, these changes lead to the formation and annihilation of SF path, creating a more complex phase transition process.

### v3

@@ III.2 Reentrance between superfluid and Bose glass phases
In the previous section, we discussed the system at relatively low filling factors. From the particle density distribution in Fig. 1(b), we observe that the average number of particles per lattice site is less than 1, indicating the influence of the on-site interaction is very weak. This thus raises the question of whether different physical phenomena might arise when the system’s filling factor is higher. To explore this, we present the results for a system with a filling factor $\rho=0.08$, and the corresponding results are summarized in Fig. 2, which includes the phase diagram and typical real-space distributions. Comparing the results to the phase diagram at low filling factors, we observe a reentrant transition at higher filling factors, as shown in Fig. 2(a). Focusing on the dashed line in this phase diagram, we see that as the on-site interaction $U$ increases, the system first transitions from the BG phase to the SF phase. However, as $U$ continues to increase, the system transitions back from the SF phase to the BG phase. This indicates that the on-site interaction can drive a reentrant transition to the BG phase.
@@
To understand this reentrant behavior, let us first check the typical real-space distributions of the system as $U$ changes shown in Fig. 2(b). When the on-site interaction $U/J=1.25$ is relatively small compared to the external potential field strength $M_{r}/J=8$, the system consists of isolated SF clusters. As we gradually increase the on-site interaction to $U/J=1.75$, a percolated SF region (marked in blue) appears in the $S_{i}$ distribution shown in Fig. 2(b), indicating that the system has transitioned from the BG phase to the SF phase.
@@
The mechanism of this stage of transition is actually similar to that of the lower-filling case presented in the previous section. Specifically, when the on-site interaction $U$ increases, the most significant interaction effect occurs at lattice sites where the average number of particles is greater than 1. At these sites, the particles repel each other, leading to a redistribution. As a result, SF paths emerge during this repulsion process, and the system transitions from the BG phase to the SF phase.
@@
However, as we further increase $U/J=3$, the system exhibits behavior different from that at low filling factors. As shown in the bottom right plot of Fig. 2(b), the percolated SF region disappears from the distribution of $S_{i}$ at $U/J=3$, indicating the system transitions from the SF phase back to the BG phase again. To understand this stage of transition, we note that although percolated SF regions can form in the SF phase, the increasing repulsive interaction strength can still cause the particle density distribution to undergo further rearrangement, causing percolated SF regions to break up into isolated SF islands, thus driving the system to enter the BG phase again. Importantly, this transition is characterized by the loss of global connectivity of the SF regions, while the local-density distribution changes only slightly. As a result, the IPR remains small [see the inset of Fig. 2(a)], even though the percolation probability $\mathcal{P}$ drops to zero, signaling the loss of long-range coherence and the onset of the BG phase. Moreover, if the particle density of each site is sufficiently low, that would indicate the local state of each site consists of only a zero- or single-occupation state. As a consequence, further increasing on-site interaction strength can hardly impose any physical influence on the system, consistent with the behavior seen in the last three columns of Fig. 1(b) for the low-filling case. In fact, at even high filling factors, one can observe even more times of reentrance (see Appendix B for more numerical results at even higher filling factors).
@@
Figure 3: (a) Phase diagram of the system at $N=100$ with quench directions marked with arrows. (b) Time dependence of the IPR of the corresponding quench dynamics. (c) Time dependence of the percolation probability $\mathcal{P}$ of the corresponding quench dynamics. Representative real-space distributions of the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, the phase $\theta_{i}\equiv\arg\langle\hat{b}_{i}\rangle$ of the SF order parameter, local winding number $w_{i}$ of the phase field $\theta_{i}$, and the discrete field $S_{i}$ during (d) intraphase and (e) interphase quench processes indicated by the arrows in (a). See text for more details.
@@ III.3 Quench dynamics
Thus far, we have focused on the static properties of the system. Now let us turn to the dynamical properties; in particular, here we shall focus on the quench dynamics of the system. Specifically, via the time-dependent Gutzwiller approach [78, 79, 80] (see Appendix C for technique details of the approach), we investigate quench dynamics of two different types, namely, intraphase quench and interphase quench.
Figure 3(d) shows the representative distributions of the system during the intraphase quenches, where the system remains within the same phase (e.g., from the BG or SF phase to the BG or SF phase). Neither the time evolution of the IPR nor the percolation probability $\mathcal{P}$ exhibits significant changes [see Figs. 3(b) and 3(c)]. Likewise, the typical real-space distributions of the amplitude of the SF order parameter and the discrete field at different time instants show no significant variations [see Fig. 3(d)]. However, we do notice the distribution of the local winding number $w_{i}$, defined for each plaquette of the lattice as the winding of the phase field $\theta_{i}\equiv\arg\langle\hat{b}_{i}\rangle$ associated with the SF order parameter, i.e., $w_{i}\equiv[(\theta_{i_{x}+1,i_{y}}-\theta_{i_{x},i_{y}})+(\theta_{i_{x}+1,i_{y}+1}-\theta_{i_{x}+1,i_{y}})+(\theta_{i_{x},i_{y}+1}-\theta_{i_{x}+1,i_{y}+1})+(\theta_{i_{x},i_{y}}-\theta_{i_{x},i_{y}+1})]/2\pi$, undergoes noticeable changes during the BG to BG quench. In this case, vortex pairs with opposite charges, indicated by the nonzero local winding number $w_{i}$, exhibit creation and annihilation dynamics. This vortex behavior can be attributed to the reorganization of the particle density distribution of the system during the quench dynamics, in which the locations of sites with vanishing particle density—fertile ground for the creation of vortex pairs—vary over time.
@@
For interphase quenches, in which the system undergoes a transition between different phases (e.g., from SF to BG or vice versa), we observe more significant qualitative changes. Specifically, for these quenches, the initial state corresponds to the ground-state configuration at the two red dots in Fig. 3(a) with $(U/J=2,M_{r}/J=5)$ in the SF phase and $(U/J=2,M_{r}/J=10)$ in the BG phase. In the quench from the SF phase to the BG phase, the percolation probability $\mathcal{P}$ [illustrated by the green line in Fig. 3(c)] rapidly decreases from nearly 1 to 0, indicating the swift transition of the system from SF to BG. This sharp drop in $\mathcal{P}$ originates from the fragmentation of the dominant percolated SF cluster into disconnected SF islands, a process driven by particle hopping in the vicinity of the percolated path; accordingly, the timescale of this drop is determined by the hopping amplitude. Once the system reaches the BG phase, $\mathcal{P}$ remains at zero, reflecting that the system remains in the BG phase. This behavior arises because the BG phase consists of isolated SF islands, which restrict particle movement to these disconnected regions. As a result, the IPR remains relatively constant with only minor fluctuations, corresponding to slight variations in the spatial distribution of particles over time [see the green line in Fig. 3(b)]. It is worth noting that the IPR does not fully rise to the value corresponding to the final BG ground state. This difference originates from the fact that the late-time state after the SF-to-BG quench is a nonequilibrium state evolving from the SF ground state, whose conserved energy differs from that of the BG ground state at the same parameters. The local-winding-number distribution $w_{i}$ also exhibits rapid dynamics following the quench, characterized by an explosive generation of vortex pairs [see the right panel in Fig. 3(e)]. This behavior signifies the breakdown of phase coherence on progressively smaller length scales.
@@
For the quench from the BG phase to the SF phase, the percolation probability $\mathcal{P}$ fluctuates between zero and nonzero values for times $tJ>5$, signaling the system’s transition between the BG and SF phases [see the blue line in Fig. 3(c)]. Unlike the quench from SF to BG, in which particles are confined to isolated SF islands, in the BG-to-SF quench, particles can move freely between SF clusters, allowing the formation of SF paths. However, at certain moments, the system may lose its SF path and transition back to the BG phase. This intermittent behavior gradually fades over time, and as shown in Fig. 3, the periods during which $\mathcal{P}=0$ become shorter, eventually leading the system to settle into a steady SF phase. This trend is also reflected in the time evolution of the IPR in Fig. 3(b). Initially, in the BG phase, the particles are localized at specific lattice sites, resulting in a relatively large IPR. As the system evolves, particles diffuse through space, causing a gradual decrease in the IPR. Although the IPR may occasionally increase, it generally decreases over time, indicating that the system eventually transitions to the SF phase. A similar pattern appears in the dynamics of the local winding number distribution $w_{i}$ [see the left panel in Fig. 3(e)]. Over time, the number of vortex pairs generally decreases, indicating the system’s progressive restoration of phase coherence across increasingly larger length scales.
@@ IV CONCLUSIONS
Our work demonstrated that ultracold bosonic systems in optical lattices with an aperiodic external potential exhibit a rich interplay between disorder and interaction effects. We showed that at low filling factors, an increase in the on-site repulsive interaction triggers a delocalization process that transforms a BG phase, characterized by isolated and disconnected SF regions, into a robust SF phase with a percolated network. This transition is clearly evidenced by the evolution of the percolation probability and IPR. At higher filling factors, we uncovered a striking reentrant behavior wherein the system initially transitions from a BG to a SF phase as the interaction strength increases, but further enhancement of the interaction causes the SF network to fragment, causing the system to revert back to a BG phase. Moreover, our analysis of quench dynamics revealed markedly different transient behaviors for intraphase and interphase quenches, highlighting the sensitive dependence of the system’s dynamical evolution on the initial state and interaction parameters. These findings not only deepen our understanding of localization phenomena in interacting Bose gases but also offer valuable insights for ongoing experiments utilizing twisted bilayer and quasicrystalline optical lattice platforms, where tuning disorder and interaction effects is pivotal for exploring novel many-body phases.
@@ Acknowledgements.
This work is supported by the NKRDPC (Grant No. 2022YFA1405304), the NSFC (Grant No. 12275089, No. 12574193, and No. 11904109), the Guangdong Basic and Applied Basic Research Foundation (Grants No. 2024A1515010188, No. 2023A1515012800, and No. 2021A1515010212), the Guangdong Provincial Key Laboratory (Grant No. 2020B1212060066), and the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401002).
V Data Availability
The data that support the findings of this article are openly available [81].
Appendix A TECHNICAL DETAILS FOR CALCULATING PERCOLATION PROBABILITY FROM A DISCRETE FIELD
In this appendix, we provide technical details for calculating the percolation probability $\mathcal{P}$ from the discrete field $S_{i}$. The percolation probability $\mathcal{P}$ is defined in Ref. [60] as
@@ Appendix A TECHNICAL DETAILS FOR CALCULATING PERCOLATION PROBABILITY FROM A DISCRETE FIELD
where $N_{\textrm{span}}$ is the number of SF sites in a percolating cluster and $N_{\varphi}$ is the total number of sites with finite $\varphi$. In our numerical calculations, $N_{\textrm{span}}$ and $N_{\varphi}$ are derived from the discrete field $S_{i}$ [77], which can be constructed by the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$ and SF order-parameter distribution $\langle\hat{b}_{i}\rangle$ described in the main text. According to the definition of $S_{i}$ [see Eq. (3)], $N_{\varphi}$ corresponds to the total number of sites where $S_{i}=1$, i.e., $N_{\varphi}=\sum_{i}^{N_{\textrm{lat}}}S_{i}$. To calculate $N_{\textrm{span}}$, it is necessary to identify the percolating cluster. The criterion for percolation, as outlined in [82], is the existence of at least one SF path that spans the entire system, either from left to right or from top to bottom. Once such an SF path is identified, the SF cluster that contains this SF path is considered the percolating cluster. The total number of lattice sites in this percolating cluster is then defined as $N_{\textrm{span}}$. Clearly, as shown by the calculation of the percolation probability $\mathcal{P}$ described above, this quantity can be used to distinguish the SF phase from the BG phase. In the BG phase, there are no percolating SF clusters, which results in a percolation probability of zero. In contrast, in the SF phase, the presence of percolating SF clusters leads to a nonzero percolation probability.
@@ Appendix B NUMERICAL RESULTS AT EVEN HIGHER FILLING FACTORS
As discussed in the main text, we hypothesize that multiple reentrance phenomena could be observed in the system with an even higher particle filling. Indeed, as anticipated, this behavior is demonstrated in Fig. 4.
Figure 4: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $1024$ particles on a $51\times 51$ square lattice. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. (b) Real-space distributions of the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ of four points along the dashed line ($M_{r}/J=25$). See text for more details.
We plot the phase diagram for a particle filling of 1024, which reveals two distinct peaks. Additionally, we includ a series of real-space distribution diagrams showing the effect of varying the interaction strength $U$ on the lattice, with the changes in $U$ indicated by the horizontal lines in the phase diagram. As shown in the inset of the phase diagram, the percolation probability of the ground state exhibits multiple transitions between zero and nonzero values, corresponding to the system’s behavior as it transitions from the BG phase to the SF phase, then back to the BG phase, and finally returns to the SF phase. The particle-density-distribution diagram Fig. 4(b) further illustrates that, as the particle filling increases, the system undergoes more significant changes due to the enhanced on-site interaction strength. Moreover, these changes lead to the formation and annihilation of SF path, creating a more complex phase transition process.
Figure 5: Dependence of the spatially averaged SF order parameter $\sum_{i}|\langle\hat{b}_{i}\rangle|/N_{\mathrm{lat}}$ on the interaction strength at different filling factors. The aperiodic potential strength is fixed at $M_{r}=7$, and the lattice size is $L=51$ ($N_{\mathrm{lat}}=51\times 51=2601$).
Moreover, at integer fillings, the standard localization effect of strong interactions can also manifest in this system, driving it into a Mott insulator phase. Figure 5 shows the dependence of the spatially averaged superfluid (SF) order parameter $\sum_{i}|\langle\hat{b}_{i}\rangle|/N_{\mathrm{lat}}$, on the interaction strength $U$ for the different fillings. At integer filling $\rho=1$, the SF order parameter vanishes at sufficiently large $U/J$, clearly signaling the emergence of a Mott insulator phase.